If an employee deposits Rs. 2,000 at the end of each year into his company’s plan which pays 7% interest compounded quarterly, how much will he have in the account at the end of 5 years?
1
Expert's answer
2019-05-23T09:12:22-0400

The amount at the end of 5 years is:

"FV = 2,000\times(1 + 0.07/4)^{20} + 2,000\times(1 + 0.07/4)^{16} + 2,000\times(1 + 0.07/4)^{12} + 2,000\times(1 + 0.07/4)^8 + 2,000\times(1 + 0.07/4)^4 + 2,000\times(1 + 0.07/4)^0 = 14,373.78."